Question 2.

A. A student titrated 10ml of 0.200M NaOH against an unknown concentration of hydrocholric acid (HCl) using phenolpthalein as an indicator. 10ml of acid was required to obtain the first permanent colour change.

i. Write the net ionic equation for this titration

ii. What was the colour change at the end point

iii. Calculate in moles/hydroxide ions originally present in the titration

iv. Calculate the unknown concentration of the hydrochloric acid

B. A 0.200M solution of methanoic acid (HCOOH) has a Ka value of 1.82 x 10⁻4

i. Calculate the equilibrium concentration of this solution

ii. Calculate the pH of the solution
1
Expert's answer
2020-05-24T03:01:18-0400

A

i) H+ + OH- = H2O

ii) The indicator changed the colour from pink to transparent

iii) n(OH-) = n(NaOH) = c(NaOH)*V(NaOH) = 0.2 mol/L*0.01 L = 0.002 mol

iv) n(HCl) = n(NaOH) = 0.002 mol; c(HCl) = n(HCl)/V(HCl) = 0.002 mol/0.01 L = 0.2 mol/L


B

i) HCOOH <-> H+ +HCOO-

Ka = c(H+)*c(HCOO-)/c(HCOOH)

Concentration of H+ ions equals x, concentration of formate anions also equals x, concentration of methanoic acid equals (0.2-x)

Ka = x*x/(0.2-x) = 1.82*10-4

x = 1.18*10-2

c(HCOOH) = 0.2-x = 0.188 mol/L

ii) pH = -lgC(H+)

c(H+) = x = 1.18*10-2

pH = 1.93